OpenStudy (anonymous):

What is the exact value of the expression 11√8+ 6√12 − 5√2? Show your work. I do believe the answer is 17√2 + 12√3, but I have to go back and check it... any help would be appreciated! :)

OpenStudy (anonymous):

look for squares in factors of radicals eg sqrt8 = sqrt (4*2) = sqrt4*sqrt2 = 2 sqrt2

OpenStudy (anonymous):

sqrt12 = 3 sqrt2

OpenStudy (anonymous):

\[11\sqrt{8}=11\sqrt{4}\sqrt{2}=22\sqrt{2}\]

OpenStudy (anonymous):

ohh okay, but what about the other one?

OpenStudy (anonymous):

look for squares in factors of radicals eg sqrt8 = sqrt (4*2) = sqrt4*sqrt2 = 2 sqrt2

OpenStudy (radar):

I think you're right Renee99

OpenStudy (anonymous):

yeah? so the other guy was wrong then?

OpenStudy (anonymous):

ill go back and check it.... I did think there was something wrong...

OpenStudy (anonymous):

renee, hold on and i will check it

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

your answer is correct

OpenStudy (anonymous):

really! awesome! :)

OpenStudy (anonymous):

thanks